OpenStudy (anonymous):

Solve x2 − 8x − 20 = 0. x = 4, x = −5 x = −4, x = 5 x = 2, x = −10 x = −2, x = 10

thomaster (thomaster):

What do you get when you factor \(\large x^2 − 8x − 20\) ?

OpenStudy (agent0smith):

You can either factor it, or plug in each value and check BOTH make the equation equal zero. ie if the equation is true for both x = 4, x = −5 (when you plug x=4 in, you get 0. And when you plug x=-5 in, you get zero), then that would be your answer.

OpenStudy (anonymous):

10*-2=-20 and 10-2=8

OpenStudy (anonymous):

\[x ^{2}-\left( 10-2 \right)x-20=0\] \[x ^{2}-10x+2x-20=0,\] make factors and get the solutions

OpenStudy (anonymous):

Thank you guys! I figured it out :)

OpenStudy (anonymous):

yw